# Linear Combinations of Logarithms of $L$-functions over Function Fields at Microscopic Shifts and Beyond

### v2

@@ Abstract.
In the function field setting with a fixed characteristic, it is known [15] that the values $\log\big|L\big(\frac{1}{2},\chi_{D}\big)\big|$ as $D$ varies over monic and square-free polynomials are asymptotically Gaussian distributed on the assumption of a low lying zeros hypothesis as the degree of $D$ tends to $\infty$. For real distinct shifts $t_{j}$ all of microscopic size or all of nonmicroscopic size relative to the genus, we consider linear combinations of $\log\big|L\big(\frac{1}{2}+it_{j},\chi_{D}\big)\big|$ with real coefficients, and separately, of $\arg L\big(\frac{1}{2}+it_{j},\chi_{D}\big).$ We provide estimates for their distribution functions under the low lying zeros hypothesis. We similarly study distribution functions of linear combinations of $\log\big|L\big(\frac{1}{2}+it_{j},E\otimes\chi_{D}\big)\big|$, and separately $\arg L\big(\frac{1}{2}+it_{j},E\otimes\chi_{D}\big)$, for quadratic twists of elliptic curves $E$ with root number one as the conductor gets large. As an application of these results, we prove a central limit theorem for the fluctuation of the number of nontrivial zeros of such $L$-functions from its mean, and thus recover previous results from [16] and [17]. Correlations of such fluctuations are in harmony with the results of [6], [12] and [34] for zeros of the Riemann zeta function and for eigenangles of unitary random matrices.
@@ 1. Introduction
A natural question to ask is about the multidimensional distribution of logarithms of distinct $L$-functions. This brings about the study of the correlations of $L$-values of type
@@
It was Bourgade [6, Theorem 1.1] who examined such local statistics for the zeta function by considering the case where the evaluation points $t_{j}$ are very close to each other. More precisely, Bourgade showed that for $\epsilon_{t}\to 0$ with $\epsilon_{t}\gg 1/\log t$ as $t\to\infty$, if the shifts are such that $0\leq t_{1}<\cdots<t_{k}<c<\infty$ as functions of $t$, and also
@@
Bourgade’s proof uses Selberg’s approximation for $\frac{\zeta^{\prime}}{\zeta}(s)$ (see [6, p. 491]) and an $L^{2}$-bound for
@@
These allow him to apply Slutsky’s lemma and the Cramér–Wald device. He also used these results to observe the correlation structure in (1.10) for fluctuations of the number of zeta zeros from its expected value (see [6, Corollary 1.3]). We note that this correlation result is the number-theoretic analogue to that of Diaconis and Evans [14] in their work for fluctuations of the number of eigenangles of random unitary matrices from its expected value.
For this paper, we were motivated to study correlations of values in the form (1.1) for symplectic and orthogonal families. A generalization of Selberg’s theorem holds for logarithms of $L$-values that form a unitary family. Since real parts of logarithms of distinct $L$-functions have approximately independent distributions on the critical line, one can obtain another central limit theorem for the distribution of linear combinations of such logarithms where coefficients are real numbers. This is a mirror image of the probabilistic fact that sums of independent Gaussian random variables are also Gaussian. This more general central limit theorem was cleverly exploited in [5] to prove that linear combinations of the $L$-functions themselves that share a functional equation, where the coefficients are again real, have a full proportion of all their nontrivial zeros on the critical line assuming suitable hypotheses on the zeros of individual $L$-functions including the Riemann hypothesis. A random matrices analogue for characteristic polynomials of independent random unitary matrices was later proven in [4].
@@ 1.1. Quadratic Dirichlet $L$-functions over Function Fields
Selberg’s theorem for the family $\{L(\tfrac{1}{2},\chi_{D})\}_{D\in\mathcal{H}_{n}}$ has been studied by the second and the third authors in [15, Theorem 1.4]. They proved that as $D\in\mathcal{H}_{n}$ and $n\to\infty$,
@@
We note that this hypothesis is motivated by the one-level density estimate (7.1), as discussed in [15, Section 2.4].
@@
In [10, Theorem 2 and Eq. (1.3)], the authors investigated the nonvanishing of various families of $L$-functions in the microscopic regime via one-level density estimates. We are thus justified in studying linear combinations of $L$-functions along microscopic shifts under the low lying zeros hypothesis.
@@
On the critical line, we will prove three theorems, each corresponding to a different regime. We say that a $k$-dimensional vector $\vec{t}$ lies in a given regime only if all of its coordinates lie in that regime. The first theorem is multidimensional and conditional.
@@ 1.2. $L$-functions of Quadratic Twists of Elliptic Curves over Function Fields
Let $E$ be an elliptic curve defined over rational numbers. The distribution of the logarithm of the (nonzero) central $L$-values $L(\tfrac{1}{2},E\otimes\chi_{d})$, where $d$ varies over fundamental discriminants was studied by Radziwiłł and Soundararajan [27, Theorem 2]. They established a one-sided central limit theorem for this orthogonal family, providing evidence for the Keating-Snaith conjectures.
@@
In other words, by combining [27, Theorem 2] with [28, Theorem 1], they established that
@@
Conditionally on this kind of low lying zeros hypothesis, one can prove an analogue of Theorem 1.3 for $\log\big|\mathfrak{L}_{\vec{a},\vec{t}}\big(\tfrac{1}{2},E\otimes\chi_{D}\big)\big|$ with microscopic $\vec{t}$ and $D\in\mathcal{H}^{\Delta,+}_{n}$ by similar methods. We skip its statement to focus on unconditional results and also to avoid confusion that can be caused by having two similar hypotheses.
@@ 1.3.2. Fluctuations of the Number of Zeros
The above result recovers a result of Faifman–Rudnick [17, eq. (1.3)] for the family $\mathcal{H}_{2g+2}$ by using the explicit formula. It was also obtained independently by Faifman [16, Theorem 3 (i)] via Selberg’s technique for the symplectic family.
The following result is another application of Theorems 1.5 and 1.7. It can be seen as evidence towards repulsion of the zeros of $L(s,\chi_{D})$ (and $L(s,E\otimes\chi_{D})$) over functions fields in the symplectic (and orthogonal) family in the mesoscopic regime. The correlations appear when counting the zeros of $\mathcal{L}(u,\mathcal{F})$ that fall in distinct intervals. The result can be compared to [6, Corollary 1.3], [12, Theorem 3.1] and [34, Theorem 1 and (3)].
@@ 1.4. Random Matrix Theory Perspective
Later in [14, Theorem 6.1], Diaconis and Evans revisited Wieand’s results on the number of eigenvalues within a fixed arc and reaffirmed the correlation structure in (1.10) by using an elegant and elementary approach based on linear combinations of traces.
The statistical properties of eigenvalues of random matrices from classical compact groups, particularly the unitary, symplectic and orthogonal groups, exhibit striking similarities to those of zeros of families of $L$-functions (see [22]). In addition to the unitary family, Keating and Snaith [24] also studied the distribution of the logarithm of characteristic polynomials at the symmetry point $(\theta=0)$ for symplectic and orthogonal families of random matrices. Regarding the local eigenvalue statistics of symplectic and orthogonal matrices, Diaconis and Evans [14, Section 8] suggested, and provided a sketch of the argument, that the correlation structure (1.10) should hold. Wieand’s approach [34] does not apply in this setting because of the complicated behavior of the corresponding Toeplitz matrices. Motivated by these insights from the random matrix theory side, we were motivated to investigate the $L$-functions side and determine whether similar local statistics of zeros arise. In fact, in Corollary 1.10, we confirm that the same phenomenon holds for our symplectic and orthogonal families of $L$-functions over function fields which will be introduced in Sections 2.2 and 2.3.
@@
By Mercator series expansion and [23, Lemma 2], we obtain the following.
@@
These coincide with the means ((averages)) appearing in Theorem 1.5 for the real and imaginary parts of the logarithm of $L$-functions over function fields. From the above, we also observe that computing the mean of the logarithm of characteristic polynomials over the family $\mathrm{USp}(2N)$ amounts to computing the mean of the trace of any power of the matrix $A$. Now, on the $L$-function side, we have an analogous product expression to (1.11), which is (2.3), but it lacks a well-structured description of the zeros of $L$-functions. Therefore, we will use Dirichlet polynomial approximations of $\log L(s,\chi_{D})$ and compute their moments instead.
Now, we proceed to compute the joint mean of the characteristic polynomial at two different shifts, say $t_{1},t_{2}\in[0,\pi)$. Using [23, Lemma 2] about the mean of the product of trace of different powers of $A$, we find that
@@
For a sketch of the proof of the multidimensional Gaussian behavior of the logarithms of characteristic polynomials, we refer the reader to [14, Section 8], which uses a different method.
@@ 2.1. Set up
One easily sees that $|\mathcal{M}_{n}|=q^{n}$. By the prime polynomial theorem [29, Theorem 2.2],
@@
Here $\mathcal{L}^{*}(u,\chi_{D})$ is a polynomial of degree $2g=d(D)-1-\eta$ (see [29, Proposition 4.3]), which we write as
@@ 2.3. $L$-functions of Quadratic Twists of Elliptic Curves
Here $\epsilon_{d(D)}\in\{\pm 1\}$ is a sign depending on the parity of $d(D)$ and $\epsilon(E)$ is the root number of the elliptic curve $E$. See [3, Lemma 2.3 and Proposition 4.3] for more details.
@@
It is further known that, see [11, Proposition 4.1],
@@
From [26, Lemma 3.4], we know that
@@
From [11, Lemma 6.2], for $(C,N_{E})=1$ and any $\epsilon>0$, one has
@@ 3.1. Orthogonality of characters over the families
Notice that $(\ell_{1},N_{E})=1$. So $\widetilde{\psi}\chi_{\ell_{1}}$ is also a primitive and non-principal character modulo $N_{E}\ell_{1}$, since $\chi_{\ell_{1}}$ is always a non-principal character. We use the Lindelöf bound (see [2, Theorem 3.4]) on $|u|=q^{-1/2}$ to get

### v3

@@ Abstract.
In the function field setting with a fixed characteristic, it is known [22] that the values $\log\big|L\big(\frac{1}{2},\chi_{D}\big)\big|$ as $D$ varies over monic and square-free polynomials are asymptotically Gaussian distributed on the assumption of a low lying zeros hypothesis as the degree of $D$ tends to $\infty$. For real distinct shifts $t_{j}$ all of microscopic size or all of nonmicroscopic size relative to the genus, we consider linear combinations of $\log\big|L\big(\frac{1}{2}+it_{j},\chi_{D}\big)\big|$ with real coefficients, and separately, of $\arg L\big(\frac{1}{2}+it_{j},\chi_{D}\big).$ We provide estimates for their distribution functions under the low lying zeros hypothesis. We similarly study distribution functions of linear combinations of $\log\big|L\big(\frac{1}{2}+it_{j},E\otimes\chi_{D}\big)\big|$, and separately $\arg L\big(\frac{1}{2}+it_{j},E\otimes\chi_{D}\big)$, for quadratic twists of elliptic curves $E$ with root number one as the conductor gets large. As an application, we prove a central limit theorem for the fluctuation of the number of nontrivial zeros of such $L$-functions from its mean, and thus recover previous results from [23] and [27]. Correlations of these fluctuations are in harmony with the results of [11], [19] and [45] for zeros of the Riemann zeta function and for eigenangles of unitary random matrices.
@@ 1. Introduction
A natural object of further inquiry is the multidimensional distribution of logarithms of distinct $L$-functions. This brings about the study of the correlations of $L$-values of type
@@
It was Bourgade [11, Theorem 1.1] who examined such local statistics for the Riemann zeta function in the case where the $t_{j}$ are close to one another. More precisely, Bourgade showed that for $\epsilon_{t}\to 0$ with $\epsilon_{t}\gg 1/\log t$ as $t\to\infty$, if the shifts are such that $0\leq t_{1}<\cdots<t_{k}<c<\infty$ as functions of $t$, and also
@@
Bourgade’s proof uses Selberg’s approximation for $\frac{\zeta^{\prime}}{\zeta}(s)$ (see [11, p. 491]) and an $L^{2}$-bound for
@@
These allow him to apply Slutsky’s lemma and the Cramér–Wald device. He also used these results to observe the correlation structure in (1.10) for fluctuations of the number of zeta zeros from its expected value (see [11, Corollary 1.3]). We note that this correlation result is the number-theoretic analogue to that of Diaconis and Evans [21] in their work for fluctuations of the number of eigenangles of random unitary matrices from its expected value. Recently, extensions of Bourgade’s result to Dirichlet $L$-functions and Dedekind zeta functions of quadratic fields were proven by Hsu and Wong [29].
For this paper, we were motivated to study correlations of values in the form (1.1) for of $L$-functions from symplectic or orthogonal families. A generalization of Selberg’s theorem holds for logarithms of $L$-functions that form a unitary family. Since real parts of logarithms of distinct $L$-functions have approximately independent distributions on the critical line, one can obtain another central limit theorem for the distribution of linear combinations of such logarithms where coefficients are real numbers. This is a mirror image of the probabilistic fact that sums of independent Gaussian random variables are also Gaussian. This more general central limit theorem was cleverly exploited in [10] to prove that linear combinations of the $L$-functions themselves that share a functional equation, where the coefficients are again real, have a full proportion of all their nontrivial zeros on the critical line assuming suitable hypotheses on the zeros of individual $L$-functions including the Riemann hypothesis. A random matrices analogue for characteristic polynomials of independent random unitary matrices was later proven in [9].
@@ 1.1. Quadratic Dirichlet $L$-functions over Function Fields
Selberg’s theorem for the family $\{L(\tfrac{1}{2},\chi_{D})\}_{D\in\mathcal{H}_{n}}$ has been studied by the second and the third authors in [22, Theorem 1.4]. They proved that as $D\in\mathcal{H}_{n}$ and $n\to\infty$,
@@
We note that this hypothesis is motivated by the one-level density estimate (7.1), as discussed in [22, Section 2.4].
@@
In [15, Theorem 2 and Eq. (1.3)], the authors investigated the nonvanishing of various families of $L$-functions in the microscopic regime via one-level density estimates. We are thus justified in studying linear combinations of $L$-functions along microscopic shifts under the low lying zeros hypothesis.
@@
On the critical line, we will prove three theorems for $t$ in different regimes. We say that a $k$-dimensional vector $\vec{t}$ lies in a given regime only if all of its coordinates lie in that regime. Our next theorem is conditional.
@@ 1.2. $L$-functions of Quadratic Twists of Elliptic Curves over Function Fields
Let $E$ be an elliptic curve defined over rational numbers. The distribution of the logarithm of the (nonzero) central $L$-values $L(\tfrac{1}{2},E\otimes\chi_{d})$, where $d$ varies over fundamental discriminants was studied by Radziwiłł and Soundararajan [38, Theorem 2]. They established a one-sided central limit theorem for this orthogonal family, providing evidence for the Keating-Snaith conjectures.
@@
In other words, by combining [38, Theorem 2] with [39, Theorem 1], they established that
@@
Conditionally on this type of low-lying zeros hypothesis, one can readily establish an analogue of Theorem 1.3 for $\log\bigl|\mathfrak{L}_{\vec{a},\vec{t}}\bigl(\tfrac{1}{2},E\otimes\chi_{D}\bigr)\bigr|$ with microscopic $\vec{t}$ and $D\in\mathcal{H}^{\Delta,+}_{n},$ using the material developed for the two families. We omit the statement in order to concentrate on the unconditional results, and also to avoid potential confusion arising from the presence of two similar hypotheses.
@@ 1.3.2. Fluctuations of the Number of Zeros
The above result recovers a result of Faifman–Rudnick [27, eq. (1.3)] for the family $\mathcal{H}_{2g+2}$ by using the explicit formula. It was also obtained independently by Faifman [23, Theorem 3 (i)] via Selberg’s technique for the symplectic family.
The following result is another application of Theorems 1.5 and 1.7. It can be seen as evidence towards repulsion of the zeros of $L(s,\chi_{D})$ (and $L(s,E\otimes\chi_{D})$) over functions fields in the symplectic (and orthogonal) family in the mesoscopic regime. The correlations appear when counting the zeros of $\mathcal{L}(u,\mathcal{F})$ that fall in distinct intervals. The result can be compared to [11, Corollary 1.3], [19, Theorem 3.1] and [45, Theorem 1 and (3)].
@@
1.3.3. Future direction of research on extreme values
The preceding theorems reveal a log-correlated Gaussian field structure for the orthogonal and symplectic families of $L$-functions over function fields. The covariance structure is likely governed by the real part of the Dirichlet polynomial given in (6.1) at various shifts. For instance, due to our computations in the proof of Proposition 6.1, the covariance between $\log|L(1/2+it_{1},\mathcal{F})|$ and $\log|L(1/2+it_{2},\mathcal{F})|$ is expected to be
(1.11) $\frac{1}{2}\log\min\left\{\kappa_{\mathcal{F}},\frac{1}{|t_{1}-t_{2}|}\right\}+\frac{1}{2}\log\min\left\{\kappa_{\mathcal{F}},\frac{1}{|t_{1}+t_{2}|}\right\},$
whereas for the unitary family of Riemann zeta functions, as shown in [11], one has for $T$ is chosen uniform from the interval $t\in[T,2T]$, that
$\mathrm{Cov}\!\left(\log\big|\zeta\big(\tfrac{1}{2}+it+it_{1}\big)\big|,\;\log\big|\zeta\big(\tfrac{1}{2}+it+it_{2}\big)\big|\right)=\frac{1}{2}\log\min\left\{\log T,\;\frac{1}{|t_{1}-t_{2}|}\right\}.$
The appearance of $|t_{1}+t_{2}|$ in (1.11) reflects the symmetry induced by the conjugate-pair structure of zeros in the orthogonal and symplectic families. In contrast, Bourgade’s covariance formula for the Riemann zeta function depends solely on $|t_{1}-t_{2}|$ and such covariance structures are characteristic of log-correlated Gaussian fields.
Using this log-correlated framework for $\zeta\big(\tfrac{1}{2}+it\big)$, Arguin, Belius, Bourgade, Radziwiłł and Soundararajan initiated a rigorous investigation of a conjecture of Fyodorov, Hiary and Keating (FHK) from [25, 26] concerning the maximum of $\zeta\big(\tfrac{1}{2}+it\big)$ on random intervals of fixed length. They established the main term predicted by the conjecture [5]. The conjecture in its full form was later proven in the works of Arguin, Bourgade and Radziwiłł [6, 7].
A key conceptual ingredient in these developments is the branching random walk analogy, which arises from the log-correlated covariance structure as depicted in [5, Figure 1]. A study of the maximum of the characteristic polynomials of random unitary matrices towards the FHK conjecture was conducted in [4]. In contrast to the unitary case, for each of the orthogonal and symplectic random matrix ensembles, the corresponding branching random walk is expected to differ substantially, and may take the form of an inhomogeneous branching random walk arising from Dyson Brownian motion as described in [17, Section 1.2.1] by Cipolloni and Landon. In loc cit, they prove a result towards the FHK conjecture for the maximum of the characteristic polynomials of i.i.d. random nonHermitian matrices. This is a recent development that makes it compelling to study the FHK conjecture for the orthogonal and symplectic families of $L$-functions that we study in this paper.
In our setting, in view of [16], a natural formulation of the FHK conjecture is as follows. Let $\Delta$ be fixed. As $n\to\infty$, for $D$ ranging over either $\mathcal{H}_{n}$ or $\mathcal{H}_{n}^{\Delta}$,
$\max_{\theta\in[0,\pi)}\log\bigl|\mathcal{L}(q^{-1/2}e(\theta),\mathcal{F})\bigr|=\sqrt{\frac{2}{\beta}}\Big(\log n-\frac{3}{4}\log\log{n}+o(1)\Big),$
where $\beta$ denotes the Dyson index which is $4$ for the symplectic family, and $1$ for the orthogonal family. It is worth noting that, for the families we consider, establishing such a result for microscopic shifts likely requires the assumption of a low lying zeros hypothesis for the real part of the logarithm. For the imaginary part, at present it is not known whether Gaussian fluctuations are exhibited at microscopic shifts. This stands in contrast with the work of Najnudel [37], who obtained the leading term for the maximum as proposed by the FHK conjecture both for the real and imaginary parts of $\log\zeta\big(\tfrac{1}{2}+it\big)$.
1.3.4. Towards ratio conjecture
The ratio conjecture for the symplectic family of quadratic Dirichlet L-functions associated with hyperelliptic curves was formulated in [2, Conjecture 6]. Our main theorem allows one to directly obtain a Selberg-type central limit theorem for the ratio of these L-functions. This, in turn, provides a useful framework for studying the ratio conjecture, since the analysis requires upper bounds for the large deviations associated with the central limit theorem (see [43]). In particular, our result gives a heuristic prediction for the order of magnitude appearing in the conjecture using (1.6) and (1.7), in accordance with the discussion on page 2 of [43].
@@ 1.4. Random Matrix Theory Perspective
Later in [21, Theorem 6.1], Diaconis and Evans revisited Wieand’s results on the number of eigenvalues within a fixed arc and reaffirmed the correlation structure in (1.10) by using an elegant and elementary approach based on linear combinations of traces.
The statistical properties of eigenvalues of random matrices from classical compact groups, particularly the unitary, symplectic and orthogonal groups, exhibit striking similarities to those of zeros of families of $L$-functions (see [32]). In addition to the unitary family, Keating and Snaith [34] also studied the distribution of the logarithm of characteristic polynomials at the symmetry point $(\theta=0)$ for symplectic and orthogonal families of random matrices. Regarding the local eigenvalue statistics of symplectic and orthogonal matrices, Diaconis and Evans [21, Section 8] suggested, and provided a sketch of the argument, that the correlation structure (1.10) should hold. Wieand’s approach [45] does not apply in this setting because of the complicated behavior of the corresponding Toeplitz matrices. Motivated by these insights from the random matrix theory side, we were motivated to investigate the $L$-functions side and determine whether similar local statistics of zeros arise. In fact, in Corollary 1.10, we confirm that the same phenomenon holds for our symplectic and orthogonal families of $L$-functions over function fields which will be introduced in Sections 2.2 and 2.3.
@@
By Mercator series expansion and [33, Lemma 2], we obtain the following.
@@
These coincide with the means ((averages)) appearing in Theorem 1.5 for the real and imaginary parts of the logarithm of $L$-functions over function fields. From the above, we also observe that computing the mean of the logarithm of characteristic polynomials over the family $\mathrm{USp}(2N)$ amounts to computing the mean of the trace of any power of the matrix $A$. Now, on the $L$-function side, we have an analogous product expression to (1.12), which is (2.3), but it lacks a well-structured description of the zeros of $L$-functions. So, we will use Dirichlet polynomial approximations of $\log L(s,\chi_{D})$ and compute their moments instead.
Now, we proceed to compute the joint mean of the characteristic polynomial at two different shifts, say $t_{1},t_{2}\in[0,\pi)$. Using [33, Lemma 2] about the mean of the product of trace of different powers of $A$, we find that
@@
For a sketch of the proof of the multidimensional Gaussian behavior of the logarithms of characteristic polynomials, we refer the reader to [21, Section 8], which uses a different method.
@@ 2.1. Set up
One easily sees that $|\mathcal{M}_{n}|=q^{n}$. By the prime polynomial theorem [40, Theorem 2.2],
@@
Here $\mathcal{L}^{*}(u,\chi_{D})$ is a polynomial of degree $2g=d(D)-1-\eta$ (see [40, Proposition 4.3]), which we write as
@@ 2.3. $L$-functions of Quadratic Twists of Elliptic Curves
Here $\epsilon_{d(D)}\in\{\pm 1\}$ is a sign depending on the parity of $d(D)$ and $\epsilon(E)$ is the root number of the elliptic curve $E$. See [8, Lemma 2.3 and Proposition 4.3] for more details.
@@
It is further known that, see [18, Proposition 4.1],
@@
From [36, Lemma 3.4], we know that
@@
From [18, Lemma 6.2], for $(C,N_{E})=1$ and any $\epsilon>0$, one has
@@ 3.1. Orthogonality of characters over the families
Notice that $(\ell_{1},N_{E})=1$. So $\widetilde{\psi}\chi_{\ell_{1}}$ is also a primitive and non-principal character modulo $N_{E}\ell_{1}$, since $\chi_{\ell_{1}}$ is always a non-principal character. We use the Lindelöf bound (see [3, Theorem 3.4]) on $|u|=q^{-1/2}$ to get